# An Introduction to Reinforcement Learning: Fundamental Concepts and Practical Applications

### v1

@@ 6.3 Off-Policy and On-Policy Methods
On-policy and off-policy learning are methodologies within model-free learning approaches, not relying on environment transition probabilities. They are classified based on the relationship between the behavior policy and the updated policy [sutton2018reinforcement].
6.3.1 On-Policy Methods
@@
On-policy methods evaluate and improve the policy used to make decisions, intertwining exploration and learning. These methods update the policy based on the actions taken and the rewards received while following the current policy. This ensures that the policy being optimized is the one actually used to interact with the environment, allowing for a coherent learning process where exploration and policy improvement are naturally integrated.
Example: Consider a customer service chatbot that learns to provide better responses to user queries. The chatbot follows a specific policy to decide which responses to give. In on-policy learning, the chatbot updates its policy based on the actual responses it uses and the feedback received from users (e.g., user satisfaction ratings). This ensures that the policy being learned is directly tied to the actions taken in real interactions, leading to stable and consistent improvement.
6.3.2 Off-Policy Methods

### v2

@@ 3.2 Off-Policy and On-Policy Methods
On-policy and off-policy learning are methodologies within model-free learning approaches, not relying on environment transition probabilities. They are classified based on the relationship between the behavior policy and the updated policy [1]. On-policy methods evaluate and improve the policy used to make decisions, intertwining exploration and learning. These methods update the policy based on the actions taken and the rewards received while following the current policy ($\pi$). This ensures that the policy being optimized is the one actually used to interact with the environment, allowing for a coherent learning process where exploration and policy improvement are naturally integrated.